Last 4 digits of CUNYFirst ID BUS 121 - - Name Time Value of Money: Problem Set la Tips: • For now, ignore payment and modKareem is wondering how many years it will take him to accumulate $50,000 for save $1,200 per year. a. How many years will it



BUS 121 - Name Last 4 digits of CUNYFirst ID Time Value of Money: Problem Set 1a Tips: . For now, ignore payment and mode
Answer #1

2(a)

Years required if saved under mattress -

$50000/$1200

= 41.67 years approx

2(b) it will take short amount of time if money is saved in an account which offers some interest because the amount of interest can be reinvested with the bank to save for making payment

3(a) janai has 44 years to save 1 million i.e. 1000000&"

Amount to save per year. = 1000000/44

=22727 approx

3(b)she will have to save less because the interest so earned can be reinvested and added to the savings

3(d)janai will have to save more of the amount needed if she decides to wait some more years after her college.

Problem set 1a

1 option-1. Borrow from bank

amount paid - 3500

Years- 3

Pv- 3000.

Rate= 3500/(3000*3)

= 0.38

Option-2 borrow from cousin

Fv- 3700

Pv- 3000

Time 4

Rate = 3700/(3000*4)

= 0.30

Answer - option 2 should be preferred as it has lower rate.

She should borrow money from her cousin